Subject: Minutes for 9/12/11 APO 3.5m users committee meeting
From: Suzanne Hawley
Submitted: Thu, 15 Sep 2011 22:18:09 -0700 (PDT)
Message number: 1246
(previous: 1245,
next: 1247
up: Index)
APO 3.5-m Users Committee Phonecon, 9/12/11
Attending: Suzanne Hawley, Mark Klaene, Bill Ketzeback, Remy Indebetouw,
Rene Walterbos (for Jon Holtzman), John Bally, Ed Turner (for Michael
Strauss), Don York (for Al Harper), Scott Anderson, Sean Moran, and
Bruce Gillespie
Absent: Russet McMillan
**********************************
Note: Three documents, relevant to today's meeting discussions, were
distributed to the Users Committee prior to the meeting:
o status of Goddard Fabry-Perot upgrade (mentioned in telescope report)
o proposal from Chris Stubbs about a new multiband imager for the 3.5-m
(the instrument formerly known as PISCO)
o document outlining the current policy for Target of Opportunity
observations on the 3.5-m
Users can request these documents from their Users Committee
representative, but the documents should not be posted in any public
location.
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - Scott had nothing to report.
o Colorado (John Bally) - John mentioned that CU is looking forward to
an upcoming student field trip to APO in March.
o Johns Hopkins (Sean Moran) - Sean had nothing to report.
o New Mexico State (Rene Walterbos) - Rene asked if DIS spectrograph focus
was checked/adjusted since the shutdown ended. He said that the spectra
he took recently seem to have 3.5 pixels fwhm. Mark said he would check.
o Princeton (Ed Turner) - Ed said he has heard nothing from users --
he did discuss the ToO topic (see below) with Michael Strauss.
o Virginia (Remy Indebetouw) - Remy had no user issues. There is a
UVa student field trip to APO planned for November.
o Chicago (Don York) - Dan has a NIC-FPS user who has been affected by
recent instrumental problems. Suzanne said that the user has been in
close contact with the site on how to deal with them. Don is working with
the site to try to understand if there have been changes in throughput at
the dichroic split. Bill said that the blue camera may have had some
contamination on the chip before the shutdown, although affecting mostly
the bluest wavelengths, and not near the dichroic split. The blue camera
was serviced during shutdown and should be fine now. He added that we
have spectrograph throughput data that can be used to check on this issue.
**********************************
Discussion of telescope/instruments report:
Mark and Bill Ketzeback reported that the telescope performance this past
month has been quite good. The replacement of the B400 grating is delayed
-- the procurement was inadvertently stuck at JHU for a while. There are
several instrument problems worth highlighting:
o the echelle slitviewer camera and controller continue to be
flaky -- both are going to be upgraded
o the slit mask wheel on TripleSpec appears to be stuck at the 1.5
arcsec slit position -- the problem is being worked with Matt Nelson (note
added 9/15: now fixed)
o NIC-FPS vacuum was unstable -- replaced ion pump, fixed (maybe)
o NIC-FPS returns corrupted images 1-2% of the frames -- under
investigation
o NIC-FPS level unbalanced between left and right sides -- may be
related to new video card installed during shutdown
Users should keep the site informed of any instrument problems they
encounter. Since the pre-shutdown problem with NIC-FPS controller resets
was thought to have been fixed during the shutdown, we are treating the
corrupted-images problem as new. John Bally will look into the
engineering support for NIC-FPS at CU.
Lastly, committee members should look at the status report on the Goddard
Fabry-Perot, which was distributed prior to the telecon. The GFP does not
yet have a TUI interface, and prospective users of the instrument are
encouraged to contact Bruce Woodgate if planning to use the instrument in
the near future.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
8/3/11 through 9/7/11
0) Overview
The monsoons continued this period, although at a significantly reduced
level. Rainfall is just over 1/2 for a normal year. Early predictions of
another La Ni?a year could mean minimal snowfall amounts this winter and
extreme forest fire danger again next year.
August saw a very busy time for the site with groups here from Goddard,
SDSS, BOSS, and APOGEE, and a student group from Univ. of Washington.
Site housing was full most of the month. We also hosted a group from the
PISCO instrument team for initial conversations on bringing the instrument
to APO.
1) Telescope
The telescope has been fully operational, and most nights this month had
no problems reported, which is a significant accomplishment following
summer shutdown. The dome appears to have taken a direct lightning strike
on August 31. The only damage was to the fire alarm system (not totally
unexpected).
Careful attention to weather by the Obs Specs has kept moisture off the
newly coated primary. A most welcome but sometimes difficult task this
time of year.
2) Instruments
DIS has settled out with most of the post-shutdown issues corrected. A
replacement Blue 400 grating is being ordered. CCD temperatures are still
running slightly elevated but closer to pre-shutdown levels.
TSpec, Agile and SPIcam operated with no new issues. Echelle is
operational. Problems with the Echelle slitviewer are still appearing,
despite a fair amount of work by site software staff. This is due to an
old computer that needs to be upgraded. We may order a new echelle
slitviewer camera soon to alleviate this issue. Additional issues have
recently surfaced with wavelength calibration and are being investigated.
The delivery of a replacement ThAr lamp with better UV throughput should
be soon. (Note added 9/15: Lamp has arrived and is now installed)
NIC-FPS began showing high ion-pump current and an occasional corrupted
image which led us to a substantial amount of vacuum work. Without
finding a smoking gun, the ion pump was replaced and the instrument cooled
down. Corrupted images continue to show up about 1% of the time.
Further troubleshooting has located ice buildup on a few pins in one of
the detector bulkhead connectors which may be the source of some of the
issues, although this connector was replaced this summer at CU. We are
currently investigating other options. On a final note, Stephane Beland,
who has been the lead CU support person for NIC-FPS for the last few
years, will be taking a new job at LASP in Boulder starting this month.
GFP commissioning characterization and software development was hampered
due to weather, but some progress was made (see attached update).
**********************************
Science Requirements for new APO imager (formerly known as PISCO):
Suzanne discussed the multi-band (griz) imager, with 12 arcminute field,
which is under construction at the CfA. A team from CfA recently came to
APO to discuss interfaces, logistics, and other topics associated with
possibly bringing the instrument here. We were left with considerable
enthusiasm for the new imager and possibilities for future instrumentation
development at Harvard. Chris Stubbs wrote a white paper describing the
instrument, which was provided to the Users Committee prior to the
meeting. We plan to develop a Science Requirements Document from this
white paper prior to holding a preliminary design review on 20-21 October.
Users Committee members should see that their users are aware of this
activity, and to give their feedback on the camera's scientific
capabilities.
John Bally asked about speciality filters for the new camera. Suzanne
said that the broad-band filters are integral to the instrument and
impossible to change. She also noted that in recent times only ~3% of the
telescope time has been used for special or narrow-band imaging. There
may be room in the new instrument where specialty filters could be mounted
-- a 3-inch filter in this position would cut the field of view to 5
arcminutes, and compromise the multi-wavelength capability in the other
passbands. We also noted that the current NA2 baffling system and guider
`limit the telescope's effective FOV to about 6 arcminutes; we would have
to remove the conical baffles to enable the full 12-arcminute field for
the instrument, which has its own internal baffling. Most other
instruments do not need the NA2 baffling, so one possibility is to remove
the existing baffling entirely. Also, there is insufficient staging space
on the NA2 side of the telescope to add another big instrument, so we
would need to remove one of our existing instruments from the staging
area.
Rene asked if DIS could be used instead for narrow-band imaging. We
thought maybe, but there are questions about the optical quality of the
imaging side of DIS. It has a ~4 arcminute FOV, and is possibly vignetted
-- this should be tested. Rene added that the NMSU 1-meter telescope is
available for general users, and it has capability for narrowbnand imaging
studies, and has a 16-arcminute FOV. Interested users should contact Rene
or Jon Holtzman.
[Note: The day following the telecon, Bill Ketzeback posted this note on
imaging with various cameras at the 3.5-m]:
"In the Users Committee Telecon a question was raised concerning the size
of narrow band imaging fields for each of the different imagers. Here is
the answer to that question.
DIS FOV 4.5 arcmin (square) 0.4 arcsec/pix unbinned
Agile FOV 2.2 arcmin (square) 0.258 arcsec/pix binned 2x2
Spicam FOV 4.7 arcmin (square) 0.282 arcsec/pix binned 2x2
GFP FOV 2.7 arcmin (diameter round) 0.16 arcsec/pix unbinned
DIS imaging of course has issues regarding which filters may be available
[i.e., usable] near the dichroic crossover. Agile has rather high read
noise for narrow band imaging."
===> ACTION: Committee members should discuss the Stubbs white paper with
their users, and solicit input for the camera's Science Requirements
Document, including types of science programs and any desired instrument
capabilities. Due by next Users Committee meeting (3 October).
**********************************
Target of Opportunity observations:
A new user at Princeton is interested in making Target of Opportunity
(ToO) observations in the fall quarter (and likely beyond). This evoked a
period in the past when we supported frequent ToO observations of GRB
detections made from the Swift satellite. We quickly learned that making
ad hoc PI-to-PI swaps to accommodate GRB follow-ups was untenable because
the PI that deferred to the ToO often felt short-changed. A policy was
then developed and used to make sure that disrupted scheduled programs
were correctly compensated -- this policy was dusted off and sent to the
Users Committee members prior to the meeting. Under this policy, ToOs and
their compensation worked fairly well, but was labor-intensive for
everyone involved. The GRB follow-up program was discontinued for various
reasons, but we came away believing that frequent ToO programs might be
best dealt with by queue scheduling blocks of time. For the current
telescope operations model that we use, ToO programs are best handled
infrequently, at the Director level. This is the practice we have been
following over the past several years.
Ed Turner said that GRB-style ToOs at a frequency of weekly is not
realistic, in his opinion, under our current operational model. But he
noted that the 3.5-m telescope was designed to make unique contributions
to ToO events, and Don York agreed. Switching to a largely
queue-scheduled operations paradigm is probably too extreme for our user
community, and it would require much more service observing (and
operational costs) and would limit the hands-on observing experience that
we feel is so valuable to our users. We then discussed compromise
options, such as giving ToO programs limited numbers of windows (possibly
pre-scheduled) per quarter. Other experiments could be considered, such
as soliciting names of PIs who would be willing to be pre-empted in
exchange for minimal compensation (eg same number of hours on a different
night, no guarantee of weather conditions). It was felt that the
follow-up of transient events may be an important capability for our
telescope in the coming years of operation, e.g., follow-up for Pan-STARRS
and LSST.
The interim solution that Suzanne adopted for the PU request was to
allocate three 2-hour blocks in the fall quarter, which may be used for
compensation time if a ToO is requested at some other time (which would
have to be arranged on a case-by-case basis). However, a long-term policy
needs to be discussed and defined.
===> ACTION: Committee members should discuss options and desiderata with
their users for a new policy for handling ToO requests on the telescope.
The summaries of these discussions should be forwarded to Suzanne for
discussion at the next Users Committee meeting.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: We will plan to phase out TUI support for MacOS 10.4 by the
end of the year, and for MacOS 10.5 by next summer. Users committee
members should query their users to see if this causes hardships.
Status: No objections to this plan were raised by the committee members.
This action is closed.
[new actions from this meeting]:
===> ACTION: Committee members should discuss the Stubbs white paper with
their users, and solicit input for the camera's Science Requirements
Document, including types of science programs and any desired instrument
capabilities. Due by next Users Committee meeting (3 October).
===> ACTION: Committee members should discuss options and desiderata with
their users for a new policy for handling ToO requests on the telescope.
The summaries of these discussions should be forwarded to Suzanne for
discussion at the next Users Committee meeting.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 3
October at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1246 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO